Solve for x.
log(x) − 2 = log(x − 7)

Answer by Alan3354(69443)

Similar problem.
========================
Solve for x.
log(x-9) - log(x) = 6
log((x-9)/x) = log(10^6)
(x-9)/x = 1000000
etc